OpenStudy (anonymous):

Simplify the expression cos(2*(arctan(x)) Really having trouble trying to simplify this thing. Would someone mind helping me out here?

myininaya (myininaya):

do you mean write as algebraic expression?

OpenStudy (anonymous):

No it just says to simplify this expression

myininaya (myininaya):

so what does that mean?

OpenStudy (anonymous):

arctanx=y tany=x required = cos(2y) = 2(cosy)^2-1 \[=2/((1+x ^{2}))-1 = (1-x ^{2})/(1+x ^{2})\]

OpenStudy (anonymous):

I am lost at the third line.

OpenStudy (anonymous):

So you basically disguise the arctan expression as a variable, and then go about with the simpliciation. ?

myininaya (myininaya):

i can write this as algebraic expression but no one is telling me what simplify means here so i don't know exactly what you want but this is what i did

OpenStudy (anonymous):

yes